1. Let Q1 = y(7), where y solves dy dx + 8x 2 = 5x, y(6) = 4. Let Q = ln(3 + |Q1|). Then T = 5 sin2 (100Q) satisfies:— (A) 0 ≤ T < 1. — (B) 1 ≤ T < 2. — (C) 2 ≤ T < 3. — (D) 3 ≤ T < 4. — (E) 4 ≤ T ≤ 5.

2. Let Q1 = y(1), where y solves dy dx + 1.7y = 5e 1.2x , y(0) = 11. Let Q = ln(3 + |Q1|). Then T = 5 sin2 (100Q) satisfies:— (A) 0 ≤ T < 1. — (B) 1 ≤ T < 2. — (C) 2 ≤ T < 3. — (D) 3 ≤ T < 4. — (E) 4 ≤ T ≤ 5.

3.10. Let Q1 = y(3), where y solves x dy dx + 4y = 11x 5 , y(2) = 11. Let Q = ln(3 + |Q1|). Then T = 5 sin2 (100Q) satisfies:— (A) 0 ≤ T < 1. — (B) 1 ≤ T < 2. — (C) 2 ≤ T < 3. — (D) 3 ≤ T < 4. — (E) 4 ≤ T ≤ 5.
Answer #1

(1 +8+ +52 : 766)=4 3 sdy = f(61-8x) de > y = 5x2 - 84 + ya)=4 → 1= @-86+ c = 4 + 8 +3672 - 57643 (= 490. 7 yfx- X+690 2 yg =dy trzy-sex Integrating factor, won solution, use the in = e dly eng - Isern um de gem - 5ler de che 7 Ce y = 5 Cem 400+41 IIntegrating factor, jüde klen - ui 2 / dy vh) = (vx xde yuh= sundon ne y = ht J- 44 Y(2)=4 → 10 c2°/-730) > - **+ 26 - m